# A Survey on Vision-Language-Action Models for Embodied AI

### v2

@@ Abstract
Deep learning has demonstrated remarkable success across many domains, including computer vision, natural language processing, and reinforcement learning. Representative artificial neural networks in these fields span convolutional neural networks, Transformers, and deep Q-networks. Built upon unimodal neural networks, numerous multi-modal models have been introduced to address a range of tasks such as visual question answering, image captioning, and speech recognition. The rise of instruction-following robotic policies in embodied AI has spurred the development of a novel category of multi-modal models known as vision-language-action models (VLAs). Their multi-modality capability has become a foundational element in robot learning. Various methods have been proposed to enhance traits such as versatility, dexterity, and generalizability. Some models focus on refining specific components. Others aim to develop control policies adept at predicting low-level actions. Certain VLAs serve as high-level task planners capable of decomposing long-horizon tasks into executable subtasks. Over the past few years, a myriad of VLAs have emerged, reflecting the rapid advancement of embodied AI. Therefore, it is imperative to capture the evolving landscape through a comprehensive survey.
@@
Embodied AI, Robotics, Multi-modality, Computer Vision, Natural Language Processing, Large Language Model
@@ I Introduction
rect/.append style=rectangle, rounded corners=2pt, dir tree switch/.style args=at #1for tree= fit=rectangle, , where level=#1 for tree= folder, grow’=0, , delay=child anchor=north, , before typesetting nodes= for tree= content/.wrap value=##1, , if=isodd(n_children(”!r”)) for nodewalk/.wrap pgfmath arg=fake=r,n=##1calign with current edgeint((n_children(”!r”)+1)/2), , , ,
{forest}
@@
Figure 1: The taxonomy of Vision-Language-Action models.
Vision-language-action models (VLAs) are a class of multi-modal models within the field of embodied AI, designed to process information from vision, language, and action modalities. Unlike conversational AI such as ChatGPT [1], embodied AI requires controlling physical embodiments that interact with the environment. Consequently, robotics is the most prominent domain of embodied AI. In language-conditioned robotic tasks, the policy must possess the capability to understand language instructions, visually perceive the environment, and generate appropriate actions, necessitating the multi-modal capabilities of VLAs. The term was recently coined by RT-2 [2]. Compared to earlier deep reinforcement learning approaches, VLAs offer superior versatility, dexterity, and generalizability in complex environments. As a result, they are well-suited not only for controlled settings like factories but also for everyday tasks in household environments [3].
Early developments in deep learning primarily consist of unimodal models. In computer vision (CV), models like AlexNet [4] showcased the potential of artificial neural networks (ANNs) [5]. Recurrent neural networks (RNNs) [6] laid the groundwork for numerous natural language processing (NLP) models but have seen a transition in recent years, with Transformers [7] taking precedence. The Deep Q-network demonstrated that ANNs can successfully tackle reinforcement learning problems. Leveraging advancements of unimodal models across diverse machine learning fields, multi-modal models have evolved to be powerful enough to tackle a variety of tasks [8], such as visual question answering, image captioning, speech recognition, etc.
Conventional robot policies based on reinforcement learning mostly concentrate on a restricted set of tasks, typically within controlled environments like factories and laboratories. For instance, [9] trains a policy specifically for grasping items. However, there is a growing need for more versatile multi-task policies, akin to recent developments in large language models (LLMs) [10, 1] and vision-language models (VLMs) [11]. Developing a multi-task policy is more challenging as it requires learning a broader set of skills and adapting to dynamic and uncertain environments. Additionally, task specification adds another layer of complexity. Some approaches use a one-hot vector to select tasks [12], but they are limited by the number of tasks in the training set.
Built upon the success of pretrained vision foundation models, LLMs, and VLMs, Vision-Language-Action models have demonstrated their competency in addressing these challenges. Pretrained visual representations from state-of-the-art vision encoders assist VLAs in perceiving complex environments [13, 14, 15], providing more precise estimations such as object class, object pose, and object geometry. With the increasing power of language models [10, 2], task specification based on language instructions becomes a viable option. Numerous ways of integrating vision models and language models have been explored by foundation VLMs, including BLIP-2 [16], Flamingo [11], etc. These innovations from different fields empower VLAs to address the challenges of embodied AI.
Different VLAs exhibit distinct emphases, as depicted by the taxonomy in Figure 1. Some works strive to enhance specific constituents of VLAs, such as pretrained visual representations, dynamics learning, and world models. Meanwhile, a substantial body of work is dedicated to robotic control policy. In this category, language instructions are input into the control policy, which generates primitive actions based on the environment. In contrast, another category of VLAs functions as high-level task planners that abstract away low-level controls. Instead, these models focus on decomposing long-horizon robotic tasks into subtasks. These subtasks can then be completed by control policies one by one, eventually fulfilling the overall task.
Related Work. Although surveys for VLAs are currently lacking, existing surveys in related domains provide valuable insights for VLA research. In the field of computer vision, surveys cover a wide range of visual models, ranging from convolutional neural networks [17] to Transformers [18]. Natural language processing models are comprehensively summarized in surveys [19, 20]. In-depth reviews of reinforcement learning can be found in surveys [21, 22, 23]. Surveys on graph neural networks are also available [24]. Furthermore, existing surveys comparing vision-language models offer inspiration for VLAs [8, 25, 26, 27]. Additionally, a survey on earlier works of embodied AI exists [28].
Contributions. This survey stands as the first thorough review of the emerging vision-language-action model in the field of embodied AI.
  • Comprehensive Review. We present a thorough review of emerging VLA models in embodied AI, covering various aspects including architectures, training objectives, and robotic tasks.
  • Taxonomy. We introduce a taxonomy of the hierarchical structure in current robot systems, comprising three main components: pretraining, control policy, and task planner. Pretraining techniques aim to enhance specific aspects of VLAs, such as the vision encoder or dynamics model. Low-level control policies execute low-level actions based on specified language commands and the perceived environment. High-level task planner decomposes long-horizon tasks into subtasks executable by control policies.
  • Abundant Resources. We provide an overview of necessary resources for training and evaluating VLA models. We investigate recently introduced datasets and simulators by comparing their key characteristics. Additionally, we include widely-adopted benchmarks for tasks such as robot control and embodied reasoning.
  • Future Directions. We outline current challenges and promising future opportunities in the field, such as addressing data scarcity, enhancing robot dexterity, achieving generalization across different tasks, environments, and embodiments, as well as improving robot safety.
Paper Organization. §II-A provides an overview of representative developments and milestones in unimodal models. Because vision-language models are closely related to vision-language-action models, recent advancements in vision-language models are compared in §II-B. §III explores various types of vision-language-action models. §IV summarizes recent datasets and benchmarks for embodied AI. Challenges and future directions are included in §V.
II Background
Figure 2: A brief timeline traces the evolution from unimodal models to multimodal models, laying the groundwork for the introduction of VLA models.
@@ II-A Unimodal Models
Vision-language-action models integrate three modalities, often relying on existing unimodal models. The transition from convolutional neural networks (e.g., ResNet [29]) to visual Transformers (e.g., ViT [30], SAM [31]) in computer vision has facilitated the development of more general vision models. In natural language processing, the evolution from recurrent neural networks (e.g., LSTM [6], GRU [32]) to Transformers [7] initially led to the pretrain-finetune paradigm (e.g., BERT [33], ChatGPT [1]), followed by the recent success of prompt tuning driven by large language models. Reinforcement learning (e.g., DQN [34], AlphaGo [35], PPO [36], Dactyl [37]) has also witnessed a shift towards employing Transformers (§III-A1) to model the Markov Decision Process as autoregressive sequential data.
@@ II-B Vision-Language Models
Vision-language tasks, encompassing image captioning [38], visual question answer [39], visual grounding [40], require the fusion of computer vision and natural language processing models. Early attempts like Show and Tell [41] capitalized on the success of early CNNs and RNNs. The introduction of high-capacity language models such as BERT [33] and GPT [42] ushered in a new era of Transformer-based VLMs, which have evolved in three major directions: self-supervised pretraining, contrastive pretraining, and multi-modal LLMs. One of the pioneering self-supervised pretraining methods is ViLBERT [43], while CLIP [44] popularized contrastive pretraining approaches for multi-modal alignment. The recent emergence of large language models has driven the development of multi-modal LLMs (MLLMs) or large multi-modal models (LMMs), which achieve state-of-the-art performance on multi-modal instruction-following tasks. Representative MLLMs include Flamingo [11], BLIP-2 [16], LLaVA [45]. VLMs share a close relationship with VLAs, as the multi-modal architectures of VLMs can be readily adopted for VLAs. For example, attaching an action decoder to VLMs can transform them into VLAs for low-level control [46]. VLMs can also serve as high-level task planners if they possess sufficient reasoning abilities.
@@ II-C Embodied AI & Robot Learning
Robot learning is also usually framed as reinforcement learning problems, represented as a Markov Decision Process (MDP) consisting of states ($s$), actions ($a$), and rewards ($r$). A MDP trajectory is denoted as $\tau=(s_{t=1},a_{t=1},r_{t=1},\dots,s_{t=T},a_{t=T},r_{t=T})$. In certain scenarios, robotic tasks may also be viewed as partially-observable Markov Decision Processes (POMDPs) due to incomplete observations. The primary objective of reinforcement learning is to train a policy capable of generating optimal actions for the current state $\pi_{\theta}(a_{t}|p,s_{\leq t},a_{<t})$. Various methods such as TD learning, policy gradients, etc., can be employed to achieve this. However, in cases where defining the reward function proves challenging, imitation learning is utilized to directly model the action distribution within trajectories devoid of rewards $\tau=(s_{t=1},a_{t=1},\dots,s_{t=T},a_{t=T})$. Furthermore, many multi-task robot models employ language as instructions to determine which task or skill to execute, thereby engaging in language-conditioned reinforcement learning.
@@ III Vision-Language-Action Models
Vision-language-action models (VLAs) are the models that handle multi-modal inputs of vision and language and output robot actions to complete embodied tasks. They serve as the cornerstone for instruction following robotic policies in the field of embodied AI. These models rely on robust vision encoders, language encoders, and action decoders. Consequently, some works focus on improving the constituent components of VLA models (Sec III-A). Others concentrate on refining low-level control policies, adept at handling simple instructions and executing primitive actions (Sec III-B). Additionally, certain VLAs abstract away from low-level control and primarily aim to solve task decomposition (Sec III-C). Thus, the combination of low-level control policy and high-level task planner can be treated as a hierarchical policy, as illustrated in Figure 3.
III-A Constituents of VLA
A variety of technologies serve as integral components of VLA models. Pioneering works such as Decision Transformer and Trajectory Transformer have paved the way for using Transformers in reinforcement learning (§III-A1), facilitating the development of scalable and high-capacity policies. The vision encoder is also crucial, as it must effectively encode the state and provide sufficient information about the environment to VLAs. Several works aim to pretrain vision encoders to obtain improved pretrained visual representations (PVRs) (§III-A2). Other research emphasizes learning the dynamics of the environment (§III-A3), enabling VLAs to understand the consequences of their actions and make more informed decisions. An extension of dynamics learning involves developing a world model (§III-A4) that can provide additional information, such as reward signals. This approach allows the model to sample imagined rollouts, reducing the need for interaction with the actual environment and improving data efficiency. Additionally, by leveraging internet-scale commonsense knowledge from large language models, LLM-induced world models (§III-A5) have been gaining popularity.
III-A1 Reinforcement Learning Transformer
Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them conducive to Transformer models. Pioneering efforts in this direction include Decision Transformer [48] and Trajectory Transformer [49]. Decision Transformer focuses on learning a policy, i.e., modeling the actions. Trajectory Transformer shares commonalities with Decision Transformer but distinguishes itself by extending its modeling objectives to include both states and returns within the trajectory. These innovative approaches showcase the applicability of Transformer architectures to the intricacies of RL sequence modeling. Gato [50] further extends this paradigm to a multi-modal, multi-task, multi-embodiment setting.
Figure 3: Illustration of a hierarchical robot policy comprising a high-level task planner and a low-level control policy. The high-level task planner generates a plan based on the user instruction, which is then executed step by step by the low-level control policy.
@@ III-A2 Pretrained Visual Representations
The effectiveness of the vision encoder directly influences the performance of the policy, as it supplies crucial information regarding object categories, positions, and environmental affordance. Consequently, numerous methods are devoted to pretraining the vision encoder to improve the quality of PVRs. Their technical details are compared in Table I.
@@
Although Contrastive Language-Image Pre-training (CLIP) [44] was not initially designed for robotics tasks, it has gained widespread adoption as a vision encoder in robotic models, such as CLIPort [13], EmbCLIP [51], and CoW [52]. The training objective of CLIP is to identify the correct text-image pair among all possible combinations in a given batch. By training to enhance alignment between the vision encoder and the language encoder, CLIP proves particularly effective in tasks where text instructions are provided as input. CLIP undergoes training on the WebImageText dataset (WIT), comprising 400 million image-text pairs. This large-scale training allows CLIP to develop a rich understanding of the relationship between visual and textual information. Notably, CLIP conducts a comprehensive comparison of various vision encoders, exploring different configurations of ResNet [29] and ViT [30]. This analysis provides valuable insights into the trade-off between accuracy and efficiency.
@@
Reusable Representations for Robotic Manipulation (R3M) [53] proposes two main pretraning objectives: time contrastive learning and video-language alignment. The objective of time contrastive learning is to minimize the distance between video frames that are temporally close while simultaneously increasing the separation between frames that are temporally distant. This objective aims to create PVRs that capture the temporal relationships within the video sequence. On the other hand, video-language alignment is to learn whether a video corresponds to a language instruction. This objective enriches the semantic relevance embedded in the PVRs.
@@
Masked Visual Pre-training (MVP) [14] adopts the methodology of Masked Autoencoder (MAE) from computer vision. The MAE involves masking out a portion of input patches to a ViT model and training it to reconstruct the corrupted patches. This approach closely resembles the masked language modeling technique used in BERT [33] and falls within the purview of self-supervised training. MVP extends this MAE objective across diverse robotic datasets, demonstrating that the pre-trained vision encoder significantly enhances performance in downstream manipulation tasks.
Value-Implicit Pre-training (VIP) [54] capitalizes on the temporal sequences present in videos, distinguishing itself from R3M [53] by attracting both the initial and target frames while simultaneously repelling successive frames. This objective aims to capture long-range temporal relationships and uphold local smoothness. While their own experiments demonstrate superior performance compared to R3M in specific tasks, subsequent research presents conflicting findings through more comprehensive evaluations [15].
Visual Cortex (VC-1) [15] conducts an in-depth examination of prior PVRs and introduces an enhanced PVR model by systematically exploring optimal ViT configurations across diverse datasets. Additionally, they perform a comprehensive comparative analysis of their model against previous methods on various manipulation and navigation datasets, shedding light on the critical factors that contribute to the improvement of PVR. Another study [55] also compares previous PVRs obtained under supervised learning or self-supervised learning.
Voltron [56] introduces a novel pretraining objective by incorporating language conditioning and language generation into the masked autoencoding (MAE) objective. Employing an encoder-decoder Transformer structure [7], the pretraining alternates between language-conditioned masked image reconstruction and language generation from masked images. This enhances the alignment between language and vision modalities, consequently resulting in a significantly improved success rate on language-conditioned imitation tasks.
RPT [57] undergoes pretraining with a focus not only on reconstructing visual inputs and robotic actions but also on proprioceptive states. In evaluating three distinct masking schemes, it was noted that token masking, in particular, yielded the most significant improvement in the model’s performance.
GR-1 [58] introduces a novel pretraining task of video prediction tailored for a GPT-style model. This video prediction objective is also subsequently employed during the finetuning phase, specifically utilizing robot data. The rationale behind this approach is that the ability to anticipate future frames contributes to more accurate action prediction. The experimental results, in the domain of robot manipulation, provide empirical support for their assertion.
SpawnNet [59] utilizes a two-stream architecture incorporating adapter layers to fuse features from both a pretrained vision encoder and features learned from scratch. This innovative approach eliminates the necessity of training the pretrained vision encoders while surpassing the performance of parameter-efficient finetuning (PEFT) methods, as evidenced by experimental results in robot manipulation tasks.
@@
Table I: Pretrained visual representations. V: Vision. L: Language. Net: Backbone network. CL: contrastive learning. MAE: masked auto-encoding. TFM: Transformer. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation. [SC]: self-collect data. For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc. $\mathcal{S}(\cdot)$ is the similarity measurement. (Ego-Data): Ego4D [60], Epic Kitchens [61], Something-Something-v2 [62], 100DOH [63].
Model V-Encoder Net Objective Formula Notation Dataset Robotic Tasks
@@
CLIP [44] ViT-B [30] - VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair WIT [SC] (EmbCLIP [51])
@@
R3M [53] ResNet-50 [29] - Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Ego4D Sim-Mani: Meta-World, Franka Kitchen, Adroit
@@
- MAE [64] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch (Ego-Data) Real-Mani (xArm7): pick, reach block, push cube, close fridge
VIP [54] ResNet-50 - Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{0},x_{j})\right)}{\exp\left(\mathcal{S}(x_{0},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{j})\right)}$ $x_{i}$: $i$-th video frame, $0<i<j$ Ego4D Real-Mani (Franka): pick and place, close drawer, push bottle, fold towel
VC-1 [15] ViT-L - MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP, including most previous datasets: (Ego-Data), Real Estate 10K [65], OpenHouse24 [SC]) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0;
Sim-Navi: Habitat
Voltron [56] TFM [7] MAP [66] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
@@
Something-Something-v2 Sim-Mani: Franka Kitchen;

### v4

@@ Abstract
Embodied AI is widely recognized as a key element of artificial general intelligence because it involves controlling embodied agents to perform tasks in the physical world. Building on the success of large language models and vision-language models, a new category of multimodal models—referred to as vision-language-action models (VLAs)—has emerged to address language-conditioned robotic tasks in embodied AI by leveraging their distinct ability to generate actions. In recent years, a myriad of VLAs have been developed, making it imperative to capture the rapidly evolving landscape through a comprehensive survey. To this end, we present the first survey on VLAs for embodied AI. This work provides a detailed taxonomy of VLAs, organized into three major lines of research. The first line focuses on individual components of VLAs. The second line is dedicated to developing control policies adept at predicting low-level actions. The third line comprises high-level task planners capable of decomposing long-horizon tasks into a sequence of subtasks, thereby guiding VLAs to follow more general user instructions. Furthermore, we provide an extensive summary of relevant resources, including datasets, simulators, and benchmarks. Finally, we discuss the challenges faced by VLAs and outline promising future directions in embodied AI. We have created a project associated with this survey, which is available at https://github.com/yueen-ma/Awesome-VLA.
@@
Vision-Language-Action Model, Embodied AI, Robotics, Multi-modality, Large Language Model, World Model
@@ I Introduction
Vision-language-action models (VLAs) are a class of multimodal models within the field of embodied AI, designed to process information from vision, language, and action modalities. Unlike conversational AI such as ChatGPT [1], embodied AI requires controlling physical embodiments that interact with the environment and robotics is the most prominent domain of embodied AI. In language-conditioned robotic tasks, the policy must possess the capability to understand language instructions, visually perceive the environment, and generate appropriate actions, necessitating the multimodal capabilities of VLAs. The term was recently coined by RT-2 [2]. Compared to earlier deep reinforcement learning approaches, VLAs offer superior versatility, dexterity, and generalizability in complex environments. As a result, they are well-suited not only for controlled settings like factories but also for everyday tasks in household environments.
Early developments in deep learning primarily consist of unimodal models. In computer vision (CV), models like AlexNet [3] showcased the potential of artifical neural networks. Recurrent neural networks laid the groundwork for numerous natural language processing (NLP) models but have seen a transition in recent years, with Transformers [4] taking precedence. The Deep Q-network [5] demonstrated that ANNs can successfully tackle reinforcement learning problems. Leveraging advancements of unimodal models across diverse machine learning fields, multimodal models now address a wide range of tasks, such as visual question answering, image captioning, and text-to-video generation.
Conventional robot policies based on reinforcement learning have largely focused on addressing a limited set of tasks within controlled environments, such as item grasping [6]. However, there is a growing need for more versatile multi-task policies, akin to recent trends in large language models (LLMs) and vision-language models (VLMs). Developing a multi-task policy is challenging, as it requires learning a broader set of skills and adapting to diverse environments. Furthermore, specifying tasks via language instructions offers a more intuitive user-robot interface, which necessitates the development of language-conditioned robot policies.
Figure 1: The general architecture of vision-language-action models. Important related components are shown in the dashed-line boxes.
Built upon the success of large VLMs, vision-language-action models have demonstrated their potential in addressing these challenges, as illustrated in Figure 1. Similar to VLMs, VLAs utilize vision foundation models as vision encoders to obtain pretrained visual representations of the current environmental state, such as object class, pose, and geometry. VLAs encode instructions using the token embeddings of their LLMs and employ various strategies to align vision and language embeddings, including approaches like BLIP-2 [7] and LLaVA [8]. By finetuning on robot data, the LLM can function as a decoder to predict actions and perform language-conditioned robotic tasks. These cross-disciplinary innovations drive the advancement of VLAs in embodied AI—a critical building block for artificial general intelligence (AGI).
(a) Concepts
(b) Timelines
Figure 2: (a) A Venn diagram that outlines the main concepts in embodied AI discussed in this paper. (b) Timelines that trace the evolution from unimodal models to vision-language-action models.
VLAs are closely related to three lines of work, as depicted by the timelines in Figure 2b and the taxonomy in Figure 3. Some approaches focus on individual components of VLAs (§III-A), such as pretrained visual representations, dynamics learning, world models, and reasoning. Meanwhile, a substantial body of research is dedicated to low-level control policies (§III-B). In this category, language instructions and visual perceptions are fed into the control policy, which then generates low-level actions—such as translation and rotation—thereby rendering VLAs an ideal choice for control policies. In contrast, another category of models serves as high-level task planners responsible for task decomposition (§IV). These models break down long-horizon tasks into a sequence of subtasks that, in turn, guide VLAs toward the overall goal, as illustrated in Figure 4. Most current robotic systems adopt such a hierarchical framework [9, 10, 11] because the high-level task planner can leverage models with high capacity while the low-level control policy can focus on speed and precision, similar to hierarchical reinforcement learning.
To provide a more comprehensive overview of current progress in embodied AI, we propose a generalized definition of “VLA,” as illustrated by the Venn diagram in Figure 2a. We define a VLA as any model capable of processing multimodal inputs from vision and language to produce robot actions that accomplish embodied tasks, typically following the architecture in Figure 1. The original concept of VLA referred to a model that adapts VLMs to robotic tasks [2]. Analogous to the distinction between LLMs and more general language models, we designate the original VLAs as “Large VLAs” (LVLAs) because they are based on LLMs or large VLMs.
Related Work. To the best of our knowledge, this survey is the first to review the recent progress of VLA models, a rapidly emerging research area. Previous surveys have investigated other facets of embodied AI. [12] comprehensively summarized foundation models in robotics up to 2023, while [13] focused on LLMs in robotics. [14] examined more recent vision, language, and robotic foundation models for general-purpose robots. [15] concentrates on real-world robot applications. In contrast, our work emphasizes VLA models, thereby complementing and extending the existing literature on embodied AI.
Contributions. To the best of our knowledge, this paper is the first comprehensive survey of emerging vision-language-action (VLA) models in the field of embodied AI.
  • Comprehensive Review. We present a thorough review of emerging VLA models in embodied AI, covering various aspects including components, architectures, training objectives, robotic tasks, etc.
  • Taxonomy. We introduce a taxonomy of VLAs based on the hierarchical framework of current robotic systems, comprising two hierarchies: a low-level control policy and a high-level task planner. Control policies execute low-level actions based on specified language commands and the perceived environment. Task planners provide guidance to control policies by decomposing long-horizon tasks into subtasks. We also discuss various essential components of VLAs.
  • Resources. We summarize the necessary resources for training and evaluating VLA models, including recent datasets and benchmarks in real-world or simulated environments. We also discuss various approaches to address current challenges, such as data scarcity and inconsistency.
  • Future Directions. We outline current challenges and promising future opportunities in the field, such as safety, foundation models, and real-world deployment.
II Background
@@
Figure 3: The taxonomy of VLA models. The organization of this survey follows this taxonomy.
@@ II-A Unimodal Models
Vision-language-action models integrate three modalities, often relying on existing unimodal models. The transition from convolutional neural networks (e.g., ResNet [16]) to visual Transformers (e.g., ViT [17], SAM [18]) in computer vision has facilitated the development of vision foundation models (VFMs). In natural language processing, the evolution from recurrent neural networks (e.g., LSTM [19], GRU [20]) to Transformers [4] initially led to the pretrain-finetune paradigm (e.g., BERT [21], ChatGPT [1]), followed by the recent success of prompt tuning driven by large language models. Reinforcement learning (e.g., DQN [5], AlphaGo [22], PPO [23], Dactyl [24]) has also witnessed a shift towards employing Transformers (§III-A1) to model the Markov Decision Process as autoregressive sequential data.
@@ II-B Vision-Language Models
Vision-language tasks, encompassing image captioning [25], visual question answering [26], visual grounding [27], require the fusion of computer vision and natural language processing models. Early efforts, such as Show and Tell [28], leveraged the success of early CNNs and RNNs. The advent of high-capacity language models, including BERT [21] and GPT [29], ushered in a new era of Transformer-based VLMs. One of the pioneering self-supervised pretraining methods is ViLBERT [30], while CLIP [31] popularized contrastive pretraining approaches for multimodal alignment. The recent emergence of large language models has driven the development of multimodal LLMs (MLLMs) or large multimodal models (LMMs), which achieve state-of-the-art performance on multimodal instruction-following tasks. Representative MLLMs include Flamingo [32], BLIP-2 [7], and LLaVA [8]. VLMs share a close relationship with VLAs, as the multimodal architectures of VLMs can be readily adopted for VLAs. Additionally, VLMs can function as high-level task planners if they possess sufficient reasoning capabilities.
@@ II-C Embodied AI & Robot Learning
Robot learning is also usually framed as reinforcement learning problems, represented as a Markov Decision Process (MDP) consisting of states ($s$), actions ($a$), and rewards ($r$). A MDP trajectory is denoted as $\tau=(s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T})$. In certain scenarios, robotic tasks may also be viewed as partially-observable Markov Decision Processes (POMDPs) due to incomplete observations. The primary objective of reinforcement learning is to train a policy capable of generating optimal actions for the current state $\pi(a_{t}|s_{\leq t},a_{<t})$. Various methods such as TD learning, policy gradients, etc., can be employed to achieve this. However, in cases where defining the reward function proves challenging, imitation learning is utilized to directly model the action distribution within trajectories devoid of rewards $\tau=(s_{1},a_{1},\dots,s_{T},a_{T})$. Furthermore, many multi-task robot models employ language as instructions $p$ to determine which task or skill to execute, leading to the development of language-conditioned robot policies $\pi(a_{t}|p,s_{\leq t},a_{<t})$.
@@ III Vision-Language-Action Models
III-A Components of VLA
A body of work focuses on individual components of VLA models, drawing on successes in CV, NLP, and RL. We introduce them through the lens of embodied AI.
III-A1 Reinforcement Learning
Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them well-suited for Transformer models. These RL Transformers have become foundational to many recent VLA models. Pioneering efforts in this direction include Decision Transformer (DT) [34] and Trajectory Transformer (TT) [35]. Gato [36] further extends this paradigm to a multi-modal/task/embodiment setting.
Reinforcement learning from human feedback (RLHF) has become an important training ingredient for LLMs. SEED [37] utilizes RLHF alongside skill-based reinforcement learning to address the sparse reward issue of long-horizon tasks in robot learning. Reflexion [38] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
@@ III-A2 Pretrained Visual Representations
The effectiveness of the vision encoder directly influences the performance of VLAs, as it provides crucial information regarding the current state $s_{t}$, such as object categories, positions, and affordance. Consequently, numerous methods are devoted to improving the quality of pretrained visual representations (PVRs). Their technical details are compared in Table I.
@@
CLIP [31] has gained widespread adoption as a vision encoder in robotic models. The training objective of CLIP is to identify the correct text-image pair among all possible combinations in a given batch. CLIP undergoes training on the WIT dataset, comprising 400 million image-text pairs. This large-scale training allows CLIP to develop a rich understanding of the relationship between visual and textual information.
@@
R3M [39] proposes two main pretraining objectives: time contrastive learning and video-language alignment. The objective of time contrastive learning is to minimize the distance between video frames that are temporally close while simultaneously increasing the separation between frames that are temporally distant. This objective aims to create PVRs that capture the temporal relationship within the video sequence. On the other hand, video-language alignment is to learn whether a video corresponds to a language instruction. This objective enriches the semantic relevance embedded in the PVRs. VIP [40] also capitalizes on video temporal relationship.
@@
MVP [41] adopts masked autoencoder (MAE) from computer vision to robotic datasets. The MAE involves masking out a portion of input patches to a ViT model and training it to reconstruct the corrupted patches. This approach closely resembles the masked language modeling technique used in BERT [21] and falls within the purview of self-supervised training. RPT [42] undergoes pretraining with a focus not only on reconstructing visual inputs but also on robotic actions and proprioceptive states.
@@
Table I: Pretrained visual representations. V: Vision. L: Language. CL: contrastive learning. TFM: Transformer. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation. For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc. $\mathcal{S}(\cdot)$ is some similarity measurement.
Model V-Encoder Type Objective Notation Robotic Tasks
@@
CLIP [31] ViT-B [17] VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair (Used by CLIPort [43], EmbCLIP [44], and CoW [45])
@@
R3M [39] ResNet-50 [16] Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Sim-Mani: Meta-World, Franka Kitchen, Adroit
@@
MAE [46] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch Real-Mani (xArm7): pick, reach block, push cube, close fridge
VIP [40] ResNet-50 Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{0},x_{j})\right)}{\exp\left(\mathcal{S}(x_{0},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{j})\right)}$ $x_{i}$: $i$-th video frame, $0<i<j$ Real-Mani (Franka): pick and place, close drawer, push bottle, fold towel
VC-1 [47] ViT-L MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP.) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0; Sim-Navi: Habitat
Voltron [48] TFM [4] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
@@
Sim-Mani: Franka Kitchen;